OpenStudy (anonymous):

Choose the correct solution in roster form. N is the set of real numbers that are factors of 12 (1 point) {1, 2, 3, 4, 6} {1, 2, 3, 4, 6, 12} {1, 2, 3, 4, 5, 6, 8, 12} {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12}

OpenStudy (eyust707):

1 * 12 = 12 2 * 6 = 12 3 * 4 = 12

OpenStudy (eyust707):

The set should only contain the numbers on the left